# SkillWrapper: Generative Predicate Invention for Skill Abstraction

### v1

@@ 3.2 Predicate Invention
$\exists\langle s_{i},\underline{\omega}_{i},s^{\prime}_{i}\rangle,\langle s_{j},\underline{\omega}_{j},s^{\prime}_{j}\rangle\in{\mathcal{D}}\text{ s.t. }s_{i}^{\prime}\in\bm{\zeta}_{\underline{\omega}_{i}},s_{j}^{\prime}\in\bm{\zeta}_{\underline{\omega}_{j}},\text{ yet }s_{i}\in{\mathcal{I}}_{\underline{\omega}_{i}}\text{ and }s_{j}\notin{\mathcal{I}}_{\underline{\omega}_{j}}.$
@@ 4 Experiments
For all experiments in this section, we consider images as fully observable state representations, assuming that an abstract state can be inferred from an image without uncertainty. Both the initial and goal states of each problem are specified using RGB images. These images may come from diverse sources, including a top-down view of an animated game, a third-person camera observing a robot, or the robot’s own egocentric perspective. All quantitative results reported in this section are averaged over three independent runs.
@@ 4.1 Implementation of Predicates
With these two properties, we can use the VLM’s response in string form as a relational predicate, and a grounded version of the predicate can be used as a classifier. We use GPT-5 (openai_gpt5) for predicate generation and evaluation.
@@ 4.2 Simulation
For each baseline that performs operator learning, we run the learning algorithm for five iterations, with each iteration proposing and executing one skill sequence consisting of 15 steps as their interaction budgets. We then evaluate each method on the evaluation set and report the average results in Table 1, where Solved % is the percentage of the problem set that was successfully solved or where impossible tasks were correctly identified by returning an empty plan, and PB stands for planning budgets—the number of plans that were tried before solving the problem (adopted from liang2025visualpredicator). We set a planning budget cap of 10 across all problems; if the planning budget has been used up for a problem, it is considered a failure. Theoretically, PB is an adequate metric that reflects the completeness of the learned model, and the impossible problems reflect its soundness. As shown in the table, SkillWrapper significantly outperforms all baselines that have no access to privileged knowledge, and even surpasses the performance of the system predicates baseline (Sys Preds.) on hard problems. Here we present the key insights, while leaving case studies, example operators, and failure modes and analysis of SkillWrapper in Appendix F.
@@
Method Easy Hard Impossible Solved % $\uparrow$ PB $\downarrow$ Solved % $\uparrow$ PB $\downarrow$ Expert Ops. 80.0 $\pm$ 4.1 2.0 $\pm$ 0.4 56.7 $\pm$ 4.1 4.8 $\pm$ 0.8 100 $\pm$ 0.0 Sys Preds. 80.0 $\pm$ 4.1 2.6 $\pm$ 0.2 16.7 $\pm$ 8.5 8.5 $\pm$ 0.8 43.3 $\pm$ 12.5 ViLa 33.3 $\pm$ 4.7 - 1.7 $\pm$ 2.4 - 16.7 $\pm$ 9.4 Random Exp. 5.0 $\pm$ 0.0 9.6 $\pm$ 0.0 0 $\pm$ 0.0 10.0 $\pm$ 0.0 100 $\pm$ 0.0 Ours 73.3 $\pm$ 2.4 2.9 $\pm$ 0.2 38.3 $\pm$ 4.1 6.1 $\pm$ 0.9 100 $\pm$ 0.0
Compared to the expert-constructed operators, which serve as an approximate upper bound for performance without privileged simulator access, SkillWrapper demonstrates competitive accuracy while requiring only a small number of exploratory interactions. In particular, SkillWrapper exhibits strong generalization from the easy to the hard set, indicating that the invented predicates and learned operators capture meaningful abstractions rather than overfitting to specific training traces. The gap between SkillWrapper and ViLa highlights the benefit of explicitly learning a symbolic model instead of relying solely on open-loop prompting, while the poor performance of random exploration underscores the importance of guided data collection in learning effective operators. These findings suggest that SkillWrapper achieves a favorable trade-off between data efficiency and model soundness and completeness, which is crucial for scaling to larger domains.
@@ 4.3 Real Robots
Figure 6: Bimanual Kuka Scenario Results over 5 iterations with invented predicate count. As the number of predicates grows, SkillWrapper improves over the baseline methods.
@@ F.2 Result Analysis of Robotouille
Random exploration is heavily limited by inefficient data gathering. However, if we take a closer look at the operators it learns, they are of good quality, which is the benefit of sharing the same predicate invention algorithm of SkillWrapper. In fact, one frequent failure mode under this randomness is that the baseline may never execute the skills with complex preconditions successfully. For example, in the Robotouille setting, the Cook action requires the item to be on the stove and the agent’s hand to be empty. Since the same predicate invention algorithm with SkillWrapper is being used, it can only learn operators for the skills that have been successfully executed in the observed transitions. As a result, it only learns operators for skills that are usually executable, such as Pick, and it thus can only solve the simplest pick and stack tasks.
Comparing the poorly performing PDDL operators learned by SkillWrapper to Expert Operators, the model only contains one extra predicate, on_cutting_board(item), which divides the previous cluster that shares the same effect into even smaller clusters. Then, the transition data in these smaller clusters cannot support the model learning algorithm to effectively eliminate spurious preconditions. This case study points out that the balance between predicate invention and data gathering is a critical factor in the learning process: if too many predicates are invented without adequate transition data, the resulting operators could possibly contain spurious preconditions, such that it cannot generalize at all. In SkillWrapper, the balance is controlled by empirically tuning the threshold of the scoring function and the length of each skill sequence proposed. Another sub-optimality is that our algorithm does not filtered invented predicates that are semantic synonyms or antonyms to existing predicates, which increases the classification burden of the foundation model. Though these redundant predicates are usually well handled by the foundation model, a smarter prompting system could be designed to mitigate this issue and improve computation efficiency.

### v2

@@ Abstract
* Equal contribution. Work done primarily when the author was a postdoc researcher at Brown University.
@@ 3.2 Predicate Invention
$\exists\langle s_{i},\underline{\omega}_{i},s^{\prime}_{i}\rangle,\langle s_{j},\underline{\omega}_{j},s^{\prime}_{j}\rangle\in{\mathcal{D}}\text{ s.t. }s_{i}^{\prime}\in\bm{\zeta}_{\underline{\omega}_{i}},s_{j}^{\prime}\in\bm{\zeta}_{\underline{\omega}_{j}},\text{ but }s_{i}\in{\mathcal{I}}_{\underline{\omega}_{i}}\text{ and }s_{j}\notin{\mathcal{I}}_{\underline{\omega}_{j}}.$
@@ 4 Experiments
For all experiments in this section, we consider images as fully observable state representations, assuming that an abstract state can be inferred from an image without uncertainty. Both the initial and goal states of each problem are specified using RGB images. These images may come from diverse sources, including a top-down view of an animated game, a third-person camera observing a robot, or the robot’s own egocentric perspective. All quantitative results reported in this section are averaged over five independent runs for simulation experiments and three runs for real robot experiments.
@@ 4.1 Implementation of Predicates
With these two properties, we can use the VLM’s response in string form as a relational predicate, and a grounded version of the predicate can be used as a classifier. We use GPT-5 (openai_gpt5) for predicate generation and evaluation. In addition to the system performance reported in the main paper, we also conducted comprehensive studies of the component-wise reliability of the VLM in Appendix G.
@@ 4.2 Simulation
  • No Heuristic: This baseline is the same as SkillWrapper, except that skill sequences are selected randomly from the foundation model’s output without applying the heuristics.
For each baseline that performs operator learning, we run the learning algorithm for five iterations, with each iteration proposing and executing one skill sequence consisting of 15 steps as their interaction budgets. We then evaluate each method on the evaluation set and report the average results in Table 1, where Solved % is the percentage of the problem set that was successfully solved or where impossible tasks were correctly identified by returning an empty plan, and PB stands for planning budgets—the number of plans that were tried before solving the problem (adopted from liang2025visualpredicator). We set a planning budget cap of 10 across all problems; if the planning budget has been used up for a problem, it is considered a failure. Theoretically, PB is an adequate metric that reflects the completeness of the learned model, and the impossible problems reflect its soundness. As shown in the table, SkillWrapper outperforms all baselines that have no access to privileged knowledge, and even surpasses the performance of the system predicates baseline (Sys Preds.) on hard problems. Here we present the key insights, while leaving case studies, example operators, and failure modes and analysis of SkillWrapper in Appendix F.
@@
Method Easy Hard Impossible Solved % $\uparrow$ PB $\downarrow$ Solved % $\uparrow$ PB $\downarrow$ Expert Ops. 81.0 $\pm$ 3.7 1.9 $\pm$ 0.4 58.1 $\pm$ 3.9 4.2 $\pm$ 0.4 100 $\pm$ 0.0 Sys Preds. 79.0 $\pm$ 3.7 2.6 $\pm$ 0.2 22.0 $\pm$ 12.9 7.8 $\pm$ 1.3 42.0 $\pm$ 7.5 ViLa 46.0 $\pm$ 16.2 - 13.9 $\pm$ 11.6 - 20.0 $\pm$ 10.9 Random Exp. 4.0 $\pm$ 2.0 9.6 $\pm$ 0.2 0 $\pm$ 0.0 10.0 $\pm$ 0.0 100 $\pm$ 0.0 No Heuristic 76.0 $\pm$ 4.9 2.5 $\pm$ 0.9 24.0 $\pm$ 19.6 7.8 $\pm$ 1.8 80 $\pm$ 20.9 Ours 74.0 $\pm$ 3.7 2.7 $\pm$ 0.4 40.0 $\pm$ 3.2 6.3 $\pm$ 0.4 100 $\pm$ 0.0
Compared to the expert-constructed operators, which serve as an approximate upper bound for performance without privileged simulator access, SkillWrapper demonstrates competitive accuracy while requiring only a small number of exploratory interactions. In particular, SkillWrapper exhibits strong generalization from the easy to the hard set, indicating that the invented predicates and learned operators capture meaningful abstractions rather than overfitting to specific training traces. The gap between SkillWrapper and ViLa highlights the benefit of explicitly learning a symbolic model instead of relying solely on open-loop prompting, while the poor performance of Random Exploration underscores the importance of guided data collection in learning effective operators, and the unstable performance of No Heuristic in hard problems shows the importance of a sufficient predicate set and how SkillWrapper manages to reliably explore the state space and learn it. These findings suggest that SkillWrapper achieves a favorable trade-off between data efficiency and model soundness and completeness, which is crucial for scaling to larger domains.
@@ 4.3 Real Robots
Figure 6: Bimanual Kuka Scenario Results over 5 iterations with invented predicate and learned operator count. As the number of predicates and operators grows, SkillWrapper improves over the baseline methods.
@@ F.2 Result Analysis of Robotouille
Random exploration is heavily limited by inefficient data gathering. However, if we take a closer look at the operators it learns, they are of good quality, which is the benefit of sharing the same predicate invention algorithm of SkillWrapper. In fact, one frequent failure mode under this randomness is that the baseline may never execute the skills with complex preconditions successfully. For example, in the Robotouille setting, the Cook action requires the item to be on the stove and the agent’s hand to be empty. Since the same predicate invention algorithm with SkillWrapper is being used, it can only learn operators for the skills that have been successfully executed in the observed transitions. As a result, it only learns operators for skills that are usually executable, such as Pick, and it thus can only solve the simplest pick and stack tasks. Additionally, random exploration also achieves $100\%$ on Impossible, yet because the learned model is not plannable.
No Heuristic shares a similar failure mode as System Predicates—achieves good performance in Easy problems but degrades significantly in Hard. With further investigation, we found that the baseline generally invents fewer predicates than SkillWrapper, which results in occasionally missing critical ones (two out of the five total runs). This observation explains the large variances in Hard and Impossible problems, and also indicates that Easy problems could be solved even with an incomplete predicate set, which aligns with our findings in System Predicate. In turn, it supports the usefulness of the two engineered heuristics for skill sequence proposal. We believe improving the exploration strategy is a promising direction for future work. Another point we want to note is that either the solved rate or the planning budget only evaluates the learned operators from the planning outcome, while better metrics are needed for evaluating the exploration efficiency.
SkillWrapper’s failure mode is similar to the case of Random exploration. Comparing the poorly performing PDDL operators learned by SkillWrapper to Expert Operators, the model only contains one extra predicate, $\mathtt{on\_cutting\_board(item)}$, which divides the previous cluster that shares the same effect into even smaller clusters. Then, the transition data in these smaller clusters cannot support the model learning algorithm to effectively eliminate spurious preconditions. This case study points out that the balance between predicate invention and data gathering is a critical factor in the learning process: if too many predicates are invented without adequate transition data, the resulting operators could possibly contain spurious preconditions, such that it cannot generalize at all. In SkillWrapper, the balance is controlled by empirically tuning the threshold of the scoring function and the length of each skill sequence proposed. Another sub-optimality is that our algorithm does not filter invented predicates that are semantic synonyms or antonyms to existing predicates, which increases the classification burden of the foundation model. Though these redundant predicates are usually well handled by the foundation model, a smarter prompting system could be designed to mitigate this issue and improve computation efficiency.
@@ G.1 Classification Accuracy
G.2 Real-world Robustness
To evaluate the real-world robustness of the VLM, we additionally conduct experiments to investigate factors such as viewpoints, lighting conditions, or domain shifts. For each of them, we collect a held-out set of images by varying these factors. We report per-predicate accuracy, and all numbers are averaged across three individual runs.
Viewpoints.
We collect visual observation data from two viewpoints and sample five configurations for each viewpoint. From the results of Franka experiments, we observe that the classification accuracies of certain predicates are higher from the viewpoint closer to the corresponding objects: at viewpoint #1, all predicates can be perfectly classified, while predicates involving gripper or mug, such as $\mathtt{gripper\_empty},\mathtt{holding}$ and $\mathtt{mug\_full}$, are significantly lower from viewpoint #2, which is farther from the objects. In bimanual Kuka experiments, the result is mostly stable across different viewpoints and generally better than in the Franka environment, which is possibly due to fewer background distractions. Though the accuracy varies across viewpoints, its performance remains reliable as long as the full observability assumption still holds.
Table 5: Per-predicate accuracy of Franka at Viewpoint #1.
Accuracy (%) 100.0 100.0 100.0 100.0 100.0 100.0
Table 6: Per-predicate accuracy of Franka at Viewpoint #2.
Accuracy (%) 80.0 80.0 90.0 100.0 100.0 100.0
(a) $\mathtt{holding(Sponge)}=\mathtt{\textbf{T}}$
(b) $\mathtt{holding(Sponge)}=\mathtt{\textbf{F}}$
Table 7: Per-predicate accuracy of Bimanual Kuka at Viewpoint #1.
Accuracy (%) 100.0 100.0 100.0 100.0 100.0 100.0
Accuracy (%) 100.0 100.0 100.0 100.0 100.0 100.0
Table 8: Per-predicate accuracy of Bimanual Kuka at Viewpoint #1.
Accuracy (%) 100.0 100.0 100.0 100.0 100.0 100.0
Accuracy (%) 100.0 100.0 100.0 100.0 100.0 93.3
(a) $\mathtt{UtensilOnTable(Knife)}=\mathtt{\textbf{T}}$
(b) $\mathtt{UtensilOnTable(Knife)}=\mathtt{\textbf{F}}$
Lighting conditions.
We collect visual observation data under two lighting conditions and sample five configurations for each one. We find that the VLM is generally robust to different lighting conditions, except for several extremely hard ones, such as under lighting condition #2 in Bimanual Kuka, where the objects are heavily shadowed.
Table 9: Per-predicate accuracy of Franka under Lighting Condition #1.
Accuracy (%) 90.0 90.0 100.0 96.7 100.0 100.0
Table 10: Per-predicate accuracy of Franka under Lighting Condition #2.
Accuracy (%) 90.0 90.0 100.0 98.3 100.0 100.0
(a) $\mathtt{stacked\_on(Teapot,Plate)}=\mathtt{\textbf{T}}$
(b) $\mathtt{stacked\_on(Teapot,Plate)}=\mathtt{\textbf{T}}$
Table 11: Per-predicate accuracy of Bimanual Kuka under Lighting Condition #1.
Accuracy (%) 100.0 100.0 100.0 100.0 100.0 100.0
Accuracy (%) 100.0 100.0 100.0 100.0 100.0 100.0
Table 12: Per-predicate accuracy of Bimanual Kuka under Lighting Condition #2.
Accuracy (%) 100.0 80.0 80.0 100.0 100.0 73.3
Accuracy (%) 100.0 100.0 100.0 86.7 100.0 100.0
(a) $\mathtt{RightGripperEmpty()}=\mathtt{\textbf{T}}$
(b) $\mathtt{RightGripperEmpty()}=\mathtt{\textbf{F}}$
Domain shift.
SkillWrapper relies entirely on semantics to prompt the VLM, abstracting raw states into symbolic states without using visual features. The only requirement for generalizing to novel objects is that the VLM can correctly identify the object referents in the image based on their type information provided in the language prompt. To evaluate this generalization capability, we collected visual observation data (five images per environment) under domain shift by swapping objects with new instances and sampling two configurations. From this observation, we found that the only failure mode introduced by domain shifts occurs when the VLM cannot recognize an object because its visual appearance does not align with the semantics. For example, a plate that is too small might be misclassified as a saucer, leading to incorrect symbolic states.
Table 13: Per-predicate accuracy of Franka under Domain Shift.
Accuracy (%) 90.0 90.0 100.0 100.0 73.3 100.0
(a) $\mathtt{stacked\_on(Teapot,Plate)}=\mathtt{\textbf{T}}$
(b) $\mathtt{stacked\_on(Teapot,Plate)}=\mathtt{\textbf{F}}$
Table 14: Per-predicate accuracy of Bimanual Kuka under Domain Shift.
Accuracy (%) 100.0 100.0 100.0 100.0 100.0 100.0
Accuracy (%) 100.0 100.0 100.0 100.0 100.0 100.0
(a) $\mathtt{coated(Knife)}=\mathtt{\textbf{T}}$
(b) $\mathtt{coated(Knife)}=\mathtt{\textbf{T}}$
G.3 Other VLMs
We further examine the possibility of using open-source VLMs as alternatives for SkillWrapper. We choose Qwen3-VL-235B (Qwen3-VL) for comparison. To evaluate its capability, we conduct two sets of preliminary experiments: predicate classification and predicate invention.
Predicate Classification.
We collected a subset of images (five from Franka and ten from Bimanual Kuka) and evaluated the truth values of each predicate with the two models. From the result, we observed that the two models have different failure patterns, and a prominent one is that Qwen3 can reliably detect if the gripper is holding an object, except for occasional classification errors on the objects being held. In general, we found two models perform on par with each other, and thus we believe they can be used interchangeably for predicate classification.
Table 15: Per-predicate accuracy of Franka.
GPT-5 Acc. (%) 60.0 60.0 100.0 100.0 100.0 100.0
Qwen3 Acc. (%) 100.0 80.0 100.0 100.0 100.0 80.0
Table 16: Per-predicate accuracy of Bimanual Kuka.
GPT-5 Acc. (%) 100.0 100.0 100.0 100.0 100.0 100.0
Qwen3 Acc. (%) 100.0 100.0 100.0 100.0 100.0 100.0
GPT-5 Acc. (%) 100.0 100.0 100.0 100.0 100.0 100.0
Qwen3 Acc. (%) 100.0 100.0 90.0 80.0 100.0 100.0
Predicate Invention.
We qualitatively compare the performance of both models on inventing predicates by reasoning over contrastive pairs of transitions. For each environment, we curated two contrastive pairs, and each model is prompted by the same input to invent one new predicate. A predicate is considered correct if it is a semantic synonym or antonym of the target predicate. From the result, we can conclude that GPT-5 is much more reliable in reasoning over the transitions for predicate invention, and thus Qwen3 cannot be used as an alternative for this specific task. (We omitted $\mathtt{?\,robot}$ from all predicates’ arguments for simplicity.)
$\mathtt{PickUp(Bowl)}$
$\mathtt{PickUp(Teapot)}$
GPT-5
$\mathtt{gripper\_is\_empty()}$
$\mathtt{gripper\_is\_empty()}$
$\mathtt{gripper\_is\_free()}$
Qwen3
$\mathtt{gripper\_free()}$
$\mathtt{gripper\_free()}$
$\mathtt{is\_visible(?pickupable)}$
Figure 17: Predicate Invention Case #1 in Franka. Target predicate: $\mathtt{GripperEmpty(}$
Existing predicates: $\emptyset$
(a) $\mathtt{Stack(Bowl,Plate)}$
$\mathtt{Stack(Bowl,Plate)}$
GPT-5
$\mathtt{plate\_top\_empty(?plate)}$
$\mathtt{plate\_is\_clean(?plate)}$
$\mathtt{plate\_is\_clean(?plate)}$
Qwen3
$\mathtt{stacked\_on}$
$\mathtt{(?pickupable,?plate)}$
$\mathtt{on\_center\_of}$
$\mathtt{(?pickupable,?plate)}$
$\mathtt{is\_fully\_supported}$
$\mathtt{(?pickupable,?plate)}$
Figure 18: Predicate Invention Case #2 in Franka. Target predicate: $\mathtt{PlateIsDirty(?\,plate)}$
Existing predicates: $\mathtt{GripperEmpty()},\mathtt{Holding(?\,pickupable)}$
$\mathtt{Scoop(Knife,Jar)}$
$\mathtt{Scoop(Knife,Jar)}$
GPT-5
$\mathtt{Open(?openable)}$
$\mathtt{Open(?openable)}$
$\mathtt{Open(?openable)}$
Qwen3
$\mathtt{UtensilInOpenable}$
$\mathtt{(?utensil,?openable)}$
$\mathtt{UtensilInOpening}$
$\mathtt{(?utensil,?openable)}$
$\mathtt{UtensilInOpenable}$
$\mathtt{(?utensil,?openable)}$
Figure 19: Predicate Invention Case #1 in Bi-Kuka. Target predicate: $\mathtt{LidOff(?\,openable)}$
[…]
$\mathtt{Open(Jar)}$
$\mathtt{Open(Jar)}$
GPT-5
$\mathtt{RightHandEmpty()}$
$\mathtt{RightHandEmpty()}$
$\mathtt{LidAttached(?openable)}$
Qwen3
$\mathtt{FullyEnclosedByLeftGripper}$
$\mathtt{(?openable)}$
$\mathtt{FullyEnclosedByLeftGripper}$
$\mathtt{(?openable)}$
$\mathtt{FullyEnclosedByLeftGripper}$
$\mathtt{(?openable)}$
Figure 20: Predicate Invention Case #2 in Bi-Kuka. Target predicate: $\mathtt{RightGripperEmpty()}$
[…]